You have been given the following information about the production of Usher Co., and are asked to provide the plant manager with information for a meeting with the vice president of operations. Standard Cost Card Direct materials (8 pounds at $5 per pound) $40.00 Direct labor (0.70 hours at $10) 7.00 Variable overhead (0.70 hours at $3 per hour) 2.10 Fixed overhead (0.70 hours at $7 per hour) 4.90 $54.00 The following is a variance report for the most recent period of operations. Variances Costs Total Standard Cost Price Quantity Direct materials $409,600 $8,378 F $9,300 U Direct labor 71,680 3,869 U 5,700 U How many pounds of raw materials were purchased and used during the period? Raw material pounds
Answer #1

Actual unit produced = 409600/40 = 10240 Units

Material quantity variance = (standard qty-actual quantity)Standard price

-9300 = (10240*8*5-5X)

5X = 409600+9300

X(actual quantity purchased and used) = 418900/5 = 83780 Pounds